OpenStudy (anonymous):

0=1-sin(2πt) ; 0≤t≤2 solve for t

OpenStudy (anonymous):

ouch,... t = 1/4, so that its sin(pi/2) = 1

OpenStudy (anonymous):

sin(0) = 0... not 1.

OpenStudy (anonymous):

so basically look for any values that make 2pi(t) = pi/2